Which of the following are solutions to the equation below? Check all that apply.
2x^2+18x=20
A. -10
B. -2
C. 1
D. 20
E. -2
F. -1

Answer by amarjeeth123(569)

2x^2+18x=20
x^2+9x-10=0
x^2+10x-x-10=0
x(x+10)-1(x+10)=0
(x-1)(x+10)=0
x=1 or x=-10
the solutions are A and C.